# MAFE: Multi-Agent Fair Environments for Decision-Making Systems

### v1

@@ C.1 Policy Action Analysis
C.2 Reward-Fairness Frontier in MAFEs
In this section, we analyze the F-MACEM algorithm’s performance in achieving fairness and accuracy, measured by the reward and fairness terms in Equation 4. Particularly, each reward and fairness violation is weighted uniformly, with $\alpha_{k}=\frac{\lambda}{K}$ for rewards and $\beta_{m}=\frac{1-\lambda}{M}$ for fairness violations. We then train the system using uniformly sampled values of $\lambda$ over the interval $[0,1]$ to analyze the trade-off between fairness and accuracy. To ensure uniform contribution from each component, we normalize all rewards and fairness violations to lie within the range $[0,1]$. The normalization factors for these results are provided in Table 7 of Appendix H.
Figure 7 presents the resulting Pareto frontiers, which illustrate the trade-off between accuracy and fairness. Each point on the frontier represents the average performance of a model trained with the same objective function across three different training seeds to represent relative fairness values. Both fairness measures from Equation 6 (for the Loan and Education environments) and Equation 7 (for the Healthcare environment) produce negative values, which are plotted directly since they are compatible with maximization. In the Loan and Education environments, fairness is assessed using a binary sensitive attribute, with a higher value indicating greater fairness. In contrast, the Healthcare environment evaluates fairness across four geographic regions, where a higher value also signifies greater fairness. In all plots, the highest fairness value corresponds to a value of 0.
C.3 Assessing the Benefit of Multi-Agent Learning
In this section, we perform an experiment to assess the benefits of allowing multiple agents to learn dynamic policies, using the Loan MAFE as a testbed. Specifically, we compare the performance of multi-agent learning, where all agents are allowed to learn optimal policies, against single-agent learning scenarios and a fixed policy baseline. The optimal policy, in this case, is defined as the one that maximizes the Loan MAFE’s objective function (as defined in equation 3), with uniform weighting applied to all terms in the objective.
Appendix D Common Considerations in MAFE Design
@@ D.2 Actions
For Agent $n$ with observation matrix, $o_{n,t}$, of size $m_{n,t}\times k_{n}$, an individual-level action takes the form $a_{n,t}\in\mathbb{R}^{m_{n,t}}$. In this case, Agent $n$ produces an action vector, where the $i^{th}$ element corresponds to a decision for the $i^{th}$ individual, whose feature vector is represented by the $i^{th}$ row of $o_{n,t}$. For instance, in the Healthcare MAFE, the Hospital agent could generate an action vector in which each element represents the priority rank assigned to an individual, determining their position in the queue for receiving an available hospital bed.
In contrast, a group-level action affects a subset of individuals in the entire population (subset of the rows of the observation matrix). The structure of a group-level action is $a_{n,t}\in\mathbb{R}^{f_{n}}$, where $f_{n}$ represents the number of decisions Agent $n$ must make, which affect all $m_{n,t}$ individuals. For example, in the Loan MAFE, the Debt Management Agent could output a single percentage value that determines the debt adjustment percentage applied to every customer’s payment at that time step. In this case the group is the entire customer repayment population.
D.3 Agents
@@
D.4 Sensitive Attribute
The sensitive attribute refers to the feature for which bias mitigation is necessary, as measured using the binary or $D$-ary metrics defined in Equations 6 and 7 in Section 5.3. In the Loan and Education MAFEs, the sensitive attribute is a binary feature indicating whether an individual belongs to an advantaged or disadvantaged group. In the Loan MAFE, this could represent attributes such as sex or race, both of which are protected characteristics under U.S. anti-discrimination laws in financial institutions (Federal Deposit Insurance Corporation, 2021). Similarly, in the Education MAFE, the sensitive attribute reflects whether an individual belongs to an underrepresented minority group at the university level.
In contrast, the Healthcare MAFE underscores that much of the disparity in health outcomes across demographic groups is driven by geographic location. For example, families of color—particularly Black families—are more likely to live in areas with limited access to healthcare facilities (U.S. Department of Health and Human Services, 2024). In this context, geographic location serves as the sensitive attribute, with four distinct regions, each associated with specific health outcome disparities.
D.5 Reward and Fairness Component Functions
@@ D.6 Transition Function
    Financial Failure: Entities like an insurance company, employer, or university may go bankrupt after incuring losses that lead to net negative profits or prevent them from paying employees.
@@
Figure 9: Loan MAFE Diagram
Table 1: Loan Processing MAFE Features
@@
Appendix E Loan MAFE
In this section we provide a detailed explanation of how we design the Loan MAFE introduced in Section 4 of the main paper.
Overview: A diagram illustrating the design of our Loan MAFE is provided in Figure 9. This environment simulates the loan processing pipeline of a financial institution. The agents in this system represent three main branches of the bank. The first is the Admissions Agent ($\boldsymbol{\pi}_{1}$), responsible for determining who will be approved for loans. The second is the Disbursement Agent ($\boldsymbol{\pi}_{2}$), which handles the timing of loan disbursements. The third is the Debt Management Agent ($\boldsymbol{\pi}_{3}$), which oversees loan repayment and manages defaults.
@@ Appendix E Loan MAFE
Table 2: Loan MAFE Component Indicators
@@
Reward and Disparity Component Indicators: At the end of time step $t$, the environment returns a collection of reward and disparity component indicators used for reward and fairness violation measurement. A summary of these indicators is provided in Table 2. Each agent in this environment represents a functioning part of one institution, namely, a bank which has one primary objective—maximizing profits ($P_{t}$). Thus, the total amount of money made by the bank at time step $t$ represents the primary reward returned by the environment. Two other rewards can constructed from this list of indicators to guide learning models to avoid poor local minima; namely overall admissions rates ($\frac{\sum_{t}\sum_{g}N_{A,t}^{g}}{\sum_{t}\sum_{g}N_{L,t}^{g}}$) and (negative) overall default rates ($-\frac{\sum_{t}\sum_{g}N_{F,t}^{g}}{\sum_{t}\sum_{g}N_{R,t}^{g}}$).
@@
Episode Termination: An episode in the Loan MAFE may terminate for two reasons: (1) The maximal number of time steps set by a user has been reached and (2) the bank goes bankrupt. Bankruptcy occurs if at any point during the simulation, the total amount of money lost by the bank is greater than the total amount of money it has received.
Appendix F Healthcare MAFE
In this section we provide a detailed explanation of how we design the Healthcare MAFE introduced in Section 4 of the main paper.
Overview: A diagram illustrating the design of our Healthcare MAFE is shown in Figure 10. This environment models the interactions among three core agents: an insurance company, a hospital, and a central planner. These agents collectively impact the health and insurance coverage of the population.
@@ Appendix F Healthcare MAFE
Table 3: Healthcare MAFE Features
@@
Table 4: Healthcare MAFE Component Indicators
@@
Indicators for Measuring Rewards and Fairness: At the end of time step $t$, the environment returns a collection of indicators used to measure rewards and fairness violations within the system. A summary of these indicators is provided in Table 4. These indicators can be used to construct the following set of rewards that motivate these agents in the real world: insurance profits ($P_{t}$), insured rates ($\frac{\sum_{t}\sum_{g}N_{g,t}^{I}}{\sum_{t}\sum_{g}N_{g,t}^{G}}$), (negative) incidence rates ($-\frac{\sum_{t}\sum_{g}N_{g,t}^{S}}{\sum_{t}\sum_{g}N_{g,t}^{H}}$), and (negative) mortality rates ($-\frac{\sum_{t}\sum_{g}N_{g,t}^{M}}{\sum_{t}\sum_{g}N_{g,t}^{T}}$).
@@
To determine if an individuals features improve, deteriorate, or remain unchanged we sample a uniform distribution over the range $[0,1]$ and update the features appropriately based on the segment in which the output value lands—$[0,P_{g,t}^{improve}]$, $(P_{g,t}^{improve},P_{g,t}^{improve}+U]$, or $(P_{g,t}^{improve}+U,1]$.
@@
Appendix G Education MAFE
Figure 13: Education MAFE Diagram
Overview: A diagram outlining the design of our Education MAFE is provided in Figure 13. This environment is designed to simulate the school-to-employment pipeline by modeling three key entities involved in this process: a university system, the employers of each individual, and an central planner (which functions as a central planner or government-like entity). The Central Planner Agent ($\boldsymbol{\pi}_{4}$) and Employer Agent ($\boldsymbol{\pi}_{3}$) are both modeled using a single agent. However, two separate agents are used to model distinct processes within the university: the Admissions Agent ($\boldsymbol{\pi}_{1}$), which determines which applicants are admitted or rejected, and the University Budget Allocation Agent ($\boldsymbol{\pi}_{2}$), which decides how to allocate the university’s budget across various expenses. The decisions made by these agents collectively shape the students’ future success.
@@ Appendix G Education MAFE
Table 5: Education MAFE Features
@@
Table 6: Education MAFE Component Indicators
@@
We obtain $f$ through training a regressor using the samples available in the IPUMS EDUC dataset where all IPUMS EDUC features from Table 5 are treated as the independent variables and GPA is treated as the dependent variable. We particularly use ridge regression for this task. $\gamma_{1}$ and $\gamma_{2}$ are user-specified weights that introduce the effect that student supports provided by the University Budget Allocation Agent have on improving student progress through the university. For these terms, we assume that ANNUALTUIT is normalized to be a percentage (between 0 and 1) and INMINTYPGRM is a binary valued variable.$\gamma_{0}\sim{\rm Uniform}[-\delta+C,\delta+C]$ is used to introduce stochasticity in baseline GPAs and is represented by uniform random noise over a window of length $2\delta$. $C$ centers this window and is adjusted based on the academic supports provided to as student. If an individual receives a significant scholarship or is provided an academic mentor, then $C>0$. Otherwise, $C=0$. Taken collectively, $f$ represents measures an individual’s baseline academic merits, while $V$ represents intervention adjusted uncertainty in an individual’s performance.
@@
To ensure that Equation G takes an inverted quadratic form, The parametrization of Equation G must be selected so that $\alpha_{2}>0$.
The intuition behind the design of Equation G is as follows. An increase in employee income leads to a marginal improvement in productivity, which directly benefits employer profits. This positive relationship is captured by the linear term in Equation G. On the other hand, paying an employee a higher salary also directly reduces the employer’s profits, which is modeled by the negative quadratic term in the same equation. The balance between these two effects depends on the interactions between employee salary and other factors captured by $\alpha_{2}$. The coefficients $\beta_{0},\dots,\beta_{3}$ can be adjusted to reflect the relative influence of these factors on employer profits. We set these values based on the intuition that higher education and better educational performance justify higher wages for employees, as they are likely to increase productivity. The quadratic term for experience captures the dual effects of greater experience: while more experience may enhance job performance, it could also lead to less flexibility in work habits and reduced exposure to the latest industry developments, as newer educational techniques and trends are typically acquired earlier in a career.
@@
We use the same modeling as was performed to model the effect of public investment in Section F to model the effect of tertiary investment for the Education MAFE, just with different application interpretation. Namely, in each time step, a subset of the updateable features in $\mathbf{v}_{v}$ associated with each individual in Region $g$ will improve with probability $P_{g,t}^{improve}$, remain unchanged with constant probability $U$, or deteriorate with probability $1-P_{g,t}^{improve}-U$. We treat $U$ as a user specified constant. The value of $P_{g,t}^{improve}$ is affected by the amount of the Central Planner’s budget that is used on tertiary investment in in Region $g$ at time step $t$. In particular, we model $P_{g,t}^{improve}$ as a function of the amount of the planners budget invested in the region in which this individual is located at time $t$. For constant hyperparameters $Q,R,V,\text{ and }W$, this is given by the following equation:
@@
To determine if an individuals features improve, deteriorate, or remain unchanged we sample a uniform distribution over the range $[0,1]$ and update the features appropriately based on the segment in which the output value lands—$[0,P_{g,t}^{improve}]$, $(P_{g,t}^{improve},P_{g,t}^{improve}+U]$, or $(P_{g,t}^{improve}+U,1]$.

### v2

@@ D.5 Policy Action Analysis
D.6 Policy Gradient Baselines for F-MACEM
To evaluate algorithmic performance under our MAFE setup, we examine three fairness-aware methods on the Loan MAFE: our proposed F-MACEM, which uses parameter-space sampling to optimize temporally aggregated fairness and utility objectives, and fairness-augmented variants of Multi-Agent Proximal Policy Optimization and Multi-Agent Deterministic Policy Gradient, denoted F-MAPPO and F-MADDPG, respectively.
These algorithms reflect distinct design assumptions. F-MAPPO and F-MADDPG are adapted from standard policy gradient methods and operate under the assumption that rewards are available as additive, per-time-step signals, using (Gaussian) noise for exploration. In our implementations, fairness and reward components are incorporated directly into the step-wise reward via a weighted combination:
$r_{t}=\sum_{k=0}^{K}\alpha_{k}r_{k,t}+\sum_{m=0}^{M}\beta_{m}f_{m,t}$ (5)
This formulation supports gradient-based learning by treating all objectives as decomposable over time. In contrast, F-MACEM uses a population-based evolutionary strategy that perturbs policy parameters and evaluates performance over entire trajectories. This allows it to optimize reward structures based on temporally aggregated statistics—such as ratio-after-aggregation fairness metrics used in MAFE-Loan. It also enables broader, high-level exploration by sampling from a distribution over full policy parameterizations, rather than relying on local action-space noise.
We compare each algorithm’s ability to balance fairness and utility according to the composite objective in Equation 1, where half the total weight is allocated to profit maximization and the remainder is evenly distributed among the fairness metrics listed in Table 1. Results for the full version of MAFE-Loan (denoted MAFE-Loan-F) are shown in Figure 11(a). Notably, F-MACEM achieves significantly better performance on the fairness-reward objective compared to F-MAPPO and F-MADDPG.
To better understand the root of this performance gap, we introduce a simplified variant of MAFE-Loan denoted MAFE-Loan-S, in which we reduce transition complexity while keeping the action, observation, and reward structures unchanged. This diagnostic setting tests whether the representational structure of our MAFEs is learnable in isolation from their dynamic complexity. If all algorithms perform well on MAFE-Loan-S, this suggests that the action-observation-reward interface is compatible with multiple learning paradigms and that the performance issues observed on MAFE-Loan-F stem from the difficulty of long-term planning under the more advanced environmental complexity of MAFE-Loan-F.
To construct the simplified environment MAFE-Loan-S, we retain the core elements of the full environment (MAFE-Loan-F)—including agent roles, action and observation formats, and the reward/fairness metrics used during training. However, we simplify the environment’s internal dynamics to reduce temporal and representational complexity while preserving the fundamental decision-making structure. In MAFE-Loan-F, individuals are tracked across multiple modules (e.g., admissions, disbursement, debt management) over extended time horizons, and system behavior is governed by separate rule-based and statistical models per module. In MAFE-Loan-S, these stages are collapsed into a single-step abstraction per individual. Instead of tracking behavior over dozens of time steps, outcomes are summarized in a single decision event using a unified rule-based transition model. Additionally, the three distinct agent observations are replaced by a single shared observation, population size and feature dimensionality are reduced, and episode length is shortened. Finally, a single logistic regression model is used for both admissions and default prediction, replacing the two distinct models used in the full version. These changes preserve the action-observation-reward interface while significantly reducing planning horizon and state evolution complexity.
Figure 11(b) confirms that each algorithm exhibits steady learning progress on MAFE-Loan-S, supporting the conclusion that the environment structure is learnable and that algorithm-environment compatibility plays a critical role in overall performance.
(a) Full MAFE-Loan
(b) Simplified MAFE-Loan
Figure 11: Performance comparison between F-MACEM, F-MAPPO, and F-MADDPG on MAFE-Loan-F (left) and MAFE-Loan-S (right).
Insights.
These results highlight how differences in exploration strategy and optimization structure can significantly affect learning outcomes in complex multi-agent environments. While all three algorithms share similar learning objectives, F-MACEM’s parameter-space exploration enables it to more effectively navigate the long-term dependencies and delayed credit assignment challenges present in MAFE-Loan-F. In contrast, F-MAPPO and F-MADDPG struggle in the full environment, likely due to their reliance on local, action-space noise and assumptions of step-wise reward decomposition. These findings underscore the importance of aligning algorithmic assumptions with environment complexity, particularly when optimizing fairness objectives that depend on temporally aggregated outcomes.
Rather than viewing this as a limitation of our MAFEs, we see it as a call to action for future algorithm development. Instead of simplifying environments to fit existing methods, future work should prioritize designing algorithms capable of engaging with the structural and temporal complexities inherent in real-world fairness-aware decision systems.
Appendix E Common Considerations in MAFE Design
@@ E.2 Actions
For Agent $n$ with observation matrix, $o_{n,t}$, of size $m_{n,t}\times k_{n}$, an individual-level action takes the form $a_{n,t}\in\mathbb{R}^{m_{n,t}}$. In this case, Agent $n$ produces an action vector, where the $i^{th}$ element corresponds to a decision for the $i^{th}$ individual, whose feature vector is represented by the $i^{th}$ row of $o_{n,t}$. For instance, in MAFE-Health, the Hospital agent could generate an action vector in which each element represents the priority rank assigned to an individual, determining their position in the queue for receiving an available hospital bed.
In contrast, a group-level action affects a subset of individuals in the entire population (subset of the rows of the observation matrix). The structure of a group-level action is $a_{n,t}\in\mathbb{R}^{f_{n}}$, where $f_{n}$ represents the number of decisions Agent $n$ must make, which affect all $m_{n,t}$ individuals. For example, in MAFE-Loan, the Debt Management Agent could output a single percentage value that determines the debt adjustment percentage applied to every customer’s payment at that time step. In this case the group is the entire customer repayment population.
E.3 Agents
@@
E.4 Sensitive Attribute
The sensitive attribute refers to the feature for which bias mitigation is necessary, as measured using the binary or $D$-ary metrics defined in Equations 3 and 4 in Section B.2. In MAFE-Loan and MAFE-Edu, the sensitive attribute is a binary feature indicating whether an individual belongs to an advantaged or disadvantaged group. In MAFE-Loan, this could represent attributes such as sex or race, both of which are protected characteristics under U.S. anti-discrimination laws in financial institutions (Federal Deposit Insurance Corporation, 2021). Similarly, in MAFE-Edu, the sensitive attribute reflects whether an individual belongs to an underrepresented minority group at the university level.
In contrast, MAFE-Health underscores that much of the disparity in health outcomes across demographic groups is driven by geographic location. For example, families of color—particularly Black families—are more likely to live in areas with limited access to healthcare facilities (U.S. Department of Health and Human Services, 2024). In this context, geographic location serves as the sensitive attribute, with four distinct regions, each associated with specific health outcome disparities.
E.5 Reward and Fairness Component Functions
@@ E.6 Transition Function
    Financial Failure: Entities like an insurance company, employer, or university may go bankrupt after incurring losses that lead to net negative profits or prevent them from paying employees.
@@
Figure 12: MAFE-Loan Diagram
Table 2: MAFE-Health Features
@@
Appendix F MAFE-Loan Modeling Details
In this section we provide a detailed explanation of how we design MAFE-Loan introduced in Section 4 of the main paper.
Overview: A diagram illustrating the design of MAFE-Loan is provided in Figure 12. This environment simulates the loan processing pipeline of a financial institution. The agents in this system represent three main branches of the bank. The first is the Admissions Agent ($\boldsymbol{\pi}_{1}$), responsible for determining who will be approved for loans. The second is the Disbursement Agent ($\boldsymbol{\pi}_{2}$), which handles the timing of loan disbursements. The third is the Debt Management Agent ($\boldsymbol{\pi}_{3}$), which oversees loan repayment and manages defaults.
@@ Appendix F MAFE-Loan Modeling Details
Table 3: MAFE-Loan Component Indicators
@@
Reward and Disparity Component Indicators: At the end of time step $t$, the environment returns a collection of reward and disparity component indicators used for reward and fairness violation measurement. A summary of these indicators is provided in Table 3. Each agent in this environment represents a functioning part of one institution, namely, a bank which has one primary objective—maximizing profits ($P_{t}$). Thus, the total amount of money made by the bank at time step $t$ represents the primary reward returned by the environment. Two other rewards can constructed from this list of indicators to guide learning models to avoid poor local minima; namely overall admissions rates ($\frac{\sum_{t}\sum_{g}N_{A,t}^{g}}{\sum_{t}\sum_{g}N_{L,t}^{g}}$) and (negative) overall default rates ($\frac{\sum_{t}\sum_{g}N_{F,t}^{g}}{\sum_{t}\sum_{g}N_{R,t}^{g}}$).
@@
Episode Termination: An episode in MAFE-Loan may terminate for two reasons: (1) The maximal number of time steps set by a user has been reached and (2) the bank goes bankrupt. Bankruptcy occurs if at any point during the simulation, the total amount of money lost by the bank is greater than the total amount of money it has received.
Appendix G MAFE-Health Modeling Details
Figure 13: MAFE-Health Diagram
In this section we provide a detailed explanation of how we design MAFE-Health introduced in Section 4 of the main paper.
Overview: A diagram illustrating the design of MAFE-Health is shown in Figure 13. This environment models the interactions among three core agents: an insurance company, a hospital, and a central planner. These agents collectively impact the health and insurance coverage of the population.
@@ Appendix G MAFE-Health Modeling Details
Table 4: MAFE-Health Features
@@
Table 5: MAFE-Health Component Indicators
@@
Indicators for Measuring Rewards and Fairness: At the end of time step $t$, the environment returns a collection of indicators used to measure rewards and fairness violations within the system. A summary of these indicators is provided in Table 5. These indicators can be used to construct the following set of rewards that motivate these agents in the real world: insurance profits ($P_{t}$), insured rates ($\frac{\sum_{t}\sum_{g}N_{g,t}^{I}}{\sum_{t}\sum_{g}N_{g,t}^{G}}$), (negative) incidence rates ($\frac{\sum_{t}\sum_{g}N_{g,t}^{S}}{\sum_{t}\sum_{g}N_{g,t}^{H}}$), and (negative) mortality rates ($\frac{\sum_{t}\sum_{g}N_{g,t}^{M}}{\sum_{t}\sum_{g}N_{g,t}^{T}}$).
@@
[…]
@@
Appendix H MAFE-Edu Modeling Details
Figure 16: MAFE-Edu Diagram
Overview: A diagram outlining the design of MAFE-Edu is provided in Figure 16. This environment is designed to simulate the school-to-employment pipeline by modeling three key entities involved in this process: a university system, the employers of each individual, and an central planner (which functions as a central planner or government-like entity). The Central Planner Agent ($\boldsymbol{\pi}_{4}$) and Employer Agent ($\boldsymbol{\pi}_{3}$) are both modeled using a single agent. However, two separate agents are used to model distinct processes within the university: the Admissions Agent ($\boldsymbol{\pi}_{1}$), which determines which applicants are admitted or rejected, and the University Budget Allocation Agent ($\boldsymbol{\pi}_{2}$), which decides how to allocate the university’s budget across various expenses. The decisions made by these agents collectively shape the students’ future success.
@@ Appendix H MAFE-Edu Modeling Details
Table 6: MAFE-Edu Features
@@
Table 7: MAFE-Edu Component Indicators
@@
[…]
@@
To ensure that Equation H takes an inverted quadratic form, The parametrization of Equation H must be selected so that $\alpha_{2}>0$.
The intuition behind the design of Equation H is as follows. An increase in employee income leads to a marginal improvement in productivity, which directly benefits employer profits. This positive relationship is captured by the linear term in Equation H. On the other hand, paying an employee a higher salary also directly reduces the employer’s profits, which is modeled by the negative quadratic term in the same equation. The balance between these two effects depends on the interactions between employee salary and other factors captured by $\alpha_{2}$. The coefficients $\beta_{0},\dots,\beta_{3}$ can be adjusted to reflect the relative influence of these factors on employer profits. We set these values based on the intuition that higher education and better educational performance justify higher wages for employees, as they are likely to increase productivity. The quadratic term for experience captures the dual effects of greater experience: while more experience may enhance job performance, it could also lead to less flexibility in work habits and reduced exposure to the latest industry developments, as newer educational techniques and trends are typically acquired earlier in a career.
@@
We use the same modeling as was performed to model the effect of public investment in Section G to model the effect of tertiary investment for MAFE-Edu, just with different application interpretation. Namely, in each time step, a subset of the updateable features in $\mathbf{v}_{v}$ associated with each individual in Region $g$ will improve with probability $P_{g,t}^{improve}$, remain unchanged with constant probability $U$, or deteriorate with probability $1-P_{g,t}^{improve}-U$. We treat $U$ as a user specified constant. The value of $P_{g,t}^{improve}$ is affected by the amount of the Central Planner’s budget that is used on tertiary investment in in Region $g$ at time step $t$. In particular, we model $P_{g,t}^{improve}$ as a function of the amount of the planners budget invested in the region in which this individual is located at time $t$. For constant hyperparameters $Q,R,V,\text{ and }W$, this is given by the following equation:
@@
[…]